# Localization phenomena in the random XXZ spin chain

### v1

@@ 3. Basic features of the XXZ spin chain
$\quad\leq C_{q}{\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}{({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}\left\lvert{\mathbf{u}_{D}}-{\mathbf{v}_{D}}\right\rvert_{1}}}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{D}(\mathbf{v}_{D},\mathbf{y}_{D})}$
$\quad\leq C_{q}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{\Lambda}{({\bf x},\mathbf{y})}}{\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}{({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}\left\lvert{\mathbf{u}_{D}}-{\mathbf{v}_{D}}\right\rvert_{1}}}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{D}(\mathbf{v}_{D},\mathbf{y}_{D})}$
$\quad\leq C_{q}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{\Lambda}{({\bf x},\mathbf{y})}}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\left(\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\mathrm{e}^{-c_{q-\frac{1}{2}}\left\lvert{\mathbf{u}_{D}}-{\mathbf{v}_{D}}\right\rvert_{1}}\right)\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{D}(\mathbf{v}_{D},\mathbf{y}_{D})}$
@@
Since $11\leq\left\lvert\mathbf{x}_{D}\right\rvert=\left\lvert\mathbf{y}_{D}\right\rvert\leq N-$ 1, we have
@@
Thus, for $z\in\mathds{H}_{q}$, using also Hölder’s inequality and the deterministic estimate (3.12). we get
@@
It now follows from (LABEL:eq:decena), (LABEL:FPxy), and (4.53) that, incorporating $s$ into the constants, that
@@
Let $q\ \in\frac{1}{2}\mathbb{N}$, $1\leq q$, and assume the the induction hypothesis, that is, Theorem 2.4 is proven for $q-\frac{1}{2}$. Let $\Lambda\subset\mathbb{Z}$ and $N\in\mathbb{N}$. Then
@@
Since $\mathbf{x}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)$, $\mathbf{x}$ has at most $\lceil q\rceil$ clusters of length $\leq N-1$, so a cluster can intersect at most two of the $S_{j}$’s (as $b>N-1$), hence $\mathbf{x}$ can intersect at most $2\lceil q\rceil$ of the $S_{j}$, $j=2,3,\ldots 6\lceil q\rceil$. It follows that there exists ${j}_{*}\in\left\{2,3,\ldots 6\lceil q\rceil-2\right\}$ such that
@@
$\left\lvert\widehat{G}_{z,q}^{D,D^{c}}(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\right\rvert\leq C_{q}\mathrm{e}^{-c_{q}\left\lvert\mathbf{u}^{\prime}-\mathbf{v}^{\prime}\right\rvert_{1}},$
@@
It then follows from (LABEL:eq:par+),(LABEL:eq:par+2), (4.64), and $\left\lvert D\right\rvert\leq 2r+1$, using also Hölder’s inequality that
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\left\lvert G_{z}^{\Lambda}({\bf x},{\bf y})\right\rvert^{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}s}}\right\}$ $\leq C_{q}\sum_{{\bf u}\in\mathcal{P}^{D}_{N}}\mathrm{e}^{-c_{q}\left\lvert{\bf x}-{\bf u}\right\rvert_{1}}$ (4.65)
$+C_{q}r^{C_{q}}\sum_{{\bf u}\in\mathcal{P}^{D}_{N}(\Lambda)}\sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}}\mathrm{e}^{-c_{q}d_{H}^{\Lambda}({\bf x},{\bf v})}\mathrm{e}^{-c_{q}\left\lvert{\bf u}-{\bf v}\right\rvert_{1}}.$
Since $\left\lvert{\bf x}-{\bf u}\right\rvert_{1}\geq\tfrac{1}{6\lceil q\rceil}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})-1$ for any ${\bf u}\in\mathcal{P}^{D}_{N}(\Lambda)$ by (4.57), we can bound
$\sum_{{\bf u}\in\mathcal{P}^{D}_{N}}\mathrm{e}^{-c_{q}\left\lvert{\bf x}-{\bf u}\right\rvert_{1}}\leq C_{q}e^{-\frac{c_{q}}{12\lceil q\rceil}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})}\sum_{{\bf u}\in\mathcal{P}_{N}}\mathrm{e}^{-\frac{c_{q}}{2}\left\lvert{\bf x}-{\bf u}\right\rvert_{1}}\leq C_{q}\mathrm{e}^{-c_{q}^{\prime}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})},$ (4.66)
@@
$d_{H}^{\Lambda}({\bf x},{\bf v})+\left\lvert{\bf u}-{\bf v}\right\rvert_{1}\geq d_{H}^{\Lambda}({\bf x},{\bf u})\geq\tfrac{1}{6\lceil q\rceil}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})-1$ (4.67)
@@
$\sum_{{\bf u}\in\mathcal{P}^{D}_{N}(\Lambda)}\sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}}\mathrm{e}^{-c_{q}d_{H}({\bf x},{\bf v})}\mathrm{e}^{-c_{q}\left\lvert{\bf u}-{\bf v}\right\rvert_{1}}\leq C_{q}r^{C^{\prime}_{q}}\mathrm{e}^{-c_{q}^{\prime}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})}\sum_{{\bf u}\in\mathcal{P}^{D}_{N}(\Lambda)}\sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}}\mathrm{e}^{-\frac{c_{q}}{2}d_{H}({\bf x},{\bf v})}\mathrm{e}^{-\frac{c_{q}}{2}\left\lvert{\bf u}-{\bf v}\right\rvert_{1}}$ (4.68)
@@
It remains consider the case $d_{H}^{\Lambda}({\bf x},{\bf y})\leq 6\lceil q\rceil N$. To do so we will prove
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\left\lvert G^{\Lambda}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}\mathrm{e}^{-c_{q}\left\lvert{\bf x}-{\bf y}\right\rvert_{1}}+C_{q}e^{-c_{q}N}\;\ \text{forall}\;\ N\in\mathbb{N},$ (4.69)
@@
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\left\lvert G^{\Lambda}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}\mathrm{e}^{-c_{q}\left\lvert{\bf x}-{\bf y}\right\rvert_{1}}+C_{q}e^{-c_{q}d_{H}^{\Lambda}({\bf x},{\bf y})}\leq C_{q}e^{-c_{q}d_{H}^{\Lambda}({\bf x},{\bf y})},$ (4.70)
which is (4.56).
To prove (4.69) we use a large deviation argument. For $N\in\mathbb{N}$, letting $S=[{\bf x}]^{\Lambda}_{N}$, let $\mathcal{E}^{S}_{N}$ be the event defined in (4.19), so we have (4.20), and (4.22) holds on the complimentary event $\left(\mathcal{E}_{N}^{S}\right)^{c}$.
@@
On the complimentary event $\left(\mathcal{E}_{N}^{S}\right)^{c}$ we use
@@
$\sup_{z\in\mathds{H}_{q}}{\left\lvert G^{S,S^{c}}_{z}({\bf x},{\bf y})\right\rvert^{s}}\leq C_{q}e^{-c_{q}\left\lvert{\bf x}-{\bf y}\right\rvert_{1}}.$ (4.73)
@@
$\leq C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial}(S)}e^{-c_{q}\left\lvert{\bf u}-{\bf y}\right\rvert_{1}}\left\lvert\left(R^{\Lambda}_{z}\Gamma\right)({\bf x},{\bf u})\right\rvert^{s}.$
@@
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\chi_{\left(\mathcal{E}_{N}^{S}\right)^{c}}\left\lvert\left(R^{\Lambda}_{z}\Gamma R^{S,S^{c}}_{z}\right)({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial}(S)}e^{-c_{q}\left\lvert{\bf u}-{\bf y}\right\rvert_{1}}\mathrm{e}^{-c_{q}d_{H}^{\Lambda}({\bf x},{\bf u})}.$ (4.75)
@@
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\chi_{\left(\mathcal{E}_{N}^{S}\right)^{c}}\left\lvert\left(R^{\Lambda}_{z}\Gamma R^{S,S^{c}}_{z}\right)({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}e^{-c_{q}N}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial}(S)}e^{-c_{q}\left\lvert{\bf u}-{\bf y}\right\rvert_{1}}\leq C_{q}\mathrm{e}^{-c_{q}N},$ (4.76)
where we used (B.3) (recall $\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)$) to get the last inequality.
Combining (4.72), (4.73), and (4.76) we get
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\chi_{\left(\mathcal{E}_{N}^{S}\right)^{c}}\left\lvert G^{\Lambda}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}e^{-c_{q}\left\lvert{\bf x}-{\bf y}\right\rvert_{1}}+C_{q}\mathrm{e}^{-c_{q}N}.$ (4.77)
@@
We will show that the theorem can be derived from [AW2, Theorem 4.1]. We start by reviewing the representation of the XXZ quantum spin chain Hamiltonian by a direct sum of discrete Schrödinger-like operators.
@@
Since $\mathcal{H}^{\Lambda},\boldsymbol{\Delta}^{\Lambda},\mathcal{W}^{\Lambda},V^{\Lambda}_{\omega}$ commute with the number of particles operator $\mathcal{N}^{\Lambda}$, they leaves each $\mathcal{H}_{\Lambda}^{\left(N\right)}$ invariant. Let $T_{N}^{\Lambda}$ be the restriction of $T^{\Lambda}$ to $\mathcal{H}_{\Lambda}^{\left(N\right)}=\ell^{2}(\Lambda^{\left(N\right)})$, where $T^{\Lambda}=\mathcal{H}^{\Lambda},\boldsymbol{\Delta}^{\Lambda},\mathcal{W}^{\Lambda},V^{\Lambda}_{\omega}$. We still have the decomposition given in (LABEL:bD):
@@
where $\boldsymbol{\Delta}^{\Lambda}_{N}$ is the adjacency operator on the graph $\Lambda^{\left(N\right)}$, $\mathcal{W}^{\Lambda}_{N}$ is a deterministic bounded potential, and $V^{\Lambda}_{N,\omega}$ is a random potential. In other words, $H_{N}^{\Lambda}$ is a random Schrödinger operator on $\ell^{2}(\Lambda^{\left(N\right)})$. For a fixed $N\in\mathbb{N}$, $H_{N}^{\Lambda}$ satisfies all the hypothesis of the operators studied on [AW2] except that it is a Schrödinger operator on $\ell^{2}(\Lambda^{\left(N\right)})$, not on $\ell^{2}(\Lambda^{N})$. This does not affect the analysis in [AW2], and all the results of [AW2] hold for $H_{N}^{\Lambda}$ for a fixed $N$.
@@
We can then apply [AW2, Theorem 4.1] to $H_{N}^{\Lambda}$ for all $N\in\mathbb{N}$. We obtain the conclusions of the theorem for $H_{N}^{\mathbb{Z}}$ for all $N\in\mathbb{N}$, with the constants independent of $N$ unless explicitly stated. It follows that the theorem holds as stated. ∎

### v3

@@ 3. Basic features of the XXZ spin chain
$\quad\leq C_{q}{\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}{({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{1}^{D}({\mathbf{u}_{D}},{\mathbf{v}_{D}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{D}(\mathbf{v}_{D},\mathbf{y}_{D})}$
$\leq C_{q}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{\Lambda}{({\bf x},\mathbf{y})}}{\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}{({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{1}^{D}({\mathbf{u}_{D}},{\mathbf{v}_{D}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{D}(\mathbf{v}_{D},\mathbf{y}_{D})}$
$\quad\leq C_{q}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{\Lambda}{({\bf x},\mathbf{y})}}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\left(\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{1}^{D}({\mathbf{u}_{D}},{\mathbf{v}_{D}})}\right)\mathrm{e}^{-c_{q-\frac{1}{2}}d_{H}^{D}(\mathbf{v}_{D},\mathbf{y}_{D})}$
@@
Since $1\leq\left\lvert\mathbf{x}_{D}\right\rvert=\left\lvert\mathbf{y}_{D}\right\rvert\leq N-1$, we have
@@
Thus, for $z\in\mathds{H}_{q}$, using Hölder’s inequality and the deterministic estimate (3.17). we get
@@
It now follows from (LABEL:eq:decena), (LABEL:FPxy), and (4.53), incorporating $s$ into the constants, that
@@
[…]
@@
Since $\mathbf{x}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)$, $\mathbf{x}$ has at most $\lceil q\rceil$ clusters of length $\leq N-1$, so a cluster can intersect at most two of the $S_{j}$’s (as b>N1b>N-1), hence $\mathbf{x}$ can intersect at most $2\lceil q\rceil$ of the $S_{j}$, $j=2,3,\ldots 6\lceil q\rceil$. It follows that there exists ${j}_{*}\in\left\{2,3,\ldots 6\lceil q\rceil-2\right\}$ such that
@@
$\left\lvert\widehat{G}_{z,q}^{D,D^{c}}(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\right\rvert\leq C_{q}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}(\mathbf{u}^{\prime},\mathbf{v}^{\prime})},$
@@
From (LABEL:eq:par+),(LABEL:eq:par+2), (4.64), and $\left\lvert D\right\rvert\leq 2r+1$, using also Hölder’s inequality, we get
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\left\lvert G_{z}^{\Lambda}({\bf x},{\bf y})\right\rvert^{s}\right\}$ $\leq C_{q}\sum_{{\bf u}\in\mathcal{P}^{D}_{N}}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}({\bf x},{\bf u})}$ (4.65)
$+C_{q}r^{C_{q}}\sum_{{\bf u}\in\mathcal{P}^{D}_{N}(\Lambda)}\sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}}\mathrm{e}^{-c_{q}d_{H}^{\Lambda}({\bf x},{\bf v})}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}({\bf u},{\bf v})}.$
Since $d_{1}^{\Lambda}(\mathbf{x},\mathbf{u})\geq\tfrac{1}{6\lceil q\rceil}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})-1$ for any ${\bf u}\in\mathcal{P}^{D}_{N}(\Lambda)$ by (4.57), we can bound
$\sum_{{\bf u}\in\mathcal{P}^{D}_{N}}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}(\mathbf{x},\mathbf{u})}\leq C_{q}e^{-\frac{c_{q}}{12\lceil q\rceil}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})}\sum_{{\bf u}\in\mathcal{P}_{N}}\mathrm{e}^{-\frac{c_{q}}{2}d_{1}^{\Lambda}(\mathbf{x},\mathbf{u})}\leq C_{q}\mathrm{e}^{-c_{q}^{\prime}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})},$ (4.66)
@@
$d_{H}^{\Lambda}({\bf x},{\bf v})+d_{1}^{\Lambda}(\mathbf{u},\mathbf{v})\geq d_{H}^{\Lambda}({\bf x},{\bf u})\geq\tfrac{1}{6\lceil q\rceil}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})-1$ (4.67)
@@
$\sum_{{\bf u}\in\mathcal{P}^{D}_{N}(\Lambda)}\sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}}\mathrm{e}^{-c_{q}d_{H}({\bf x},{\bf v})}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}(\mathbf{u},\mathbf{v})}\leq C_{q}r^{C^{\prime}_{q}}\mathrm{e}^{-c_{q}^{\prime}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})}\sum_{{\bf u}\in\mathcal{P}^{D}_{N}(\Lambda)}\sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}}\mathrm{e}^{-\frac{c_{q}}{2}d_{H}({\bf x},{\bf v})}\mathrm{e}^{-\frac{c_{q}}{2}d_{1}^{\Lambda}(\mathbf{u},\mathbf{v})}$ (4.68)
@@
It remains consider the case $d_{H}^{\Lambda}({\bf x},{\bf y})\leq 6\lceil q\rceil N$. To do so, we will show that
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\left\lvert G^{\Lambda}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})}+C_{q}e^{-c_{q}N}\;\ \text{for all}\;\ N\in\mathbb{N},$ (4.69)
@@
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\left\lvert G^{\Lambda}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})}+C_{q}e^{-c_{q}d_{H}^{\Lambda}({\bf x},{\bf y})}\leq C_{q}e^{-c_{q}d_{H}^{\Lambda}({\bf x},{\bf y})},$ (4.70)
establishing (4.56).
To prove (4.69) we use a large deviation argument. For $N\in\mathbb{N}$, letting $S=[{\bf x}]^{\Lambda}_{N}$, let $\mathcal{E}^{S}_{N}$ be the event defined in (4.19), so we have (4.20), and (4.22) holds on the complementary event $\left(\mathcal{E}_{N}^{S}\right)^{c}$.
@@
On the complementary event $\left(\mathcal{E}_{N}^{S}\right)^{c}$ we use
@@
$\sup_{z\in\mathds{H}_{q}}{\left\lvert G^{S,S^{c}}_{z}({\bf x},{\bf y})\right\rvert^{s}}\leq C_{q}e^{-c_{q}d_{1}^{S}(\mathbf{x},\mathbf{y})}\leq C_{q}e^{-c_{q}d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})}.$ (4.73)
@@
$\leq C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial}(S)}e^{-c_{q}d_{1}^{\Lambda}(\mathbf{u},\mathbf{y})}\left\lvert\left(R^{\Lambda}_{z}\Gamma\right)({\bf x},{\bf u})\right\rvert^{s}.$
@@
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\chi_{\left(\mathcal{E}_{N}^{S}\right)^{c}}\left\lvert\left(R^{\Lambda}_{z}\Gamma R^{S,S^{c}}_{z}\right)({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial}(S)}e^{-c_{q}d_{1}^{\Lambda}(\mathbf{u},\mathbf{y})}\mathrm{e}^{-c_{q}d_{H}^{\Lambda}({\bf x},{\bf u})}.$ (4.75)
@@
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\chi_{\left(\mathcal{E}_{N}^{S}\right)^{c}}\left\lvert\left(R^{\Lambda}_{z}\Gamma R^{S,S^{c}}_{z}\right)({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}e^{-c_{q}N}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial}(S)}e^{-c_{q}d_{1}^{\Lambda}(\mathbf{u},\mathbf{y})}\leq C_{q}\mathrm{e}^{-c_{q}N},$ (4.76)
where we used (B.3) (recall $\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)$) to get the last inequality. Combining (4.72), (4.73), and (4.76) we get
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\chi_{\left(\mathcal{E}_{N}^{S}\right)^{c}}\left\lvert G^{\Lambda}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}e^{-c_{q}d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})}+C_{q}\mathrm{e}^{-c_{q}N}.$ (4.77)
@@
We will show that the theorem can be derived from [3, Theorem 4.1]. We start by reviewing the representation of the XXZ quantum spin chain Hamiltonian by a direct sum of discrete Schrödinger-like operators.
@@
Since $H^{\Lambda},\boldsymbol{\Delta}^{\Lambda},\mathcal{W}^{\Lambda},V^{\Lambda}_{\omega}$ commute with the number of particles operator $\mathcal{N}^{\Lambda}$, they leave each $\mathcal{H}_{\Lambda}^{\left(N\right)}$ invariant. Let $T_{N}^{\Lambda}$ be the restriction of $T^{\Lambda}$ to $\mathcal{H}_{\Lambda}^{\left(N\right)}=\ell^{2}(\Lambda^{\left(N\right)})$, where $T^{\Lambda}=H^{\Lambda},\boldsymbol{\Delta}^{\Lambda},\mathcal{W}^{\Lambda},V^{\Lambda}_{\omega}$. We still have the decomposition given in (LABEL:bD):
@@
where $\boldsymbol{\Delta}^{\Lambda}_{N}$ is the adjacency operator on the graph $\Lambda^{\left(N\right)}$, $\mathcal{W}^{\Lambda}_{N}$ is a deterministic bounded potential, and $V^{\Lambda}_{N,\omega}$ is a random potential. In other words, $H_{N}^{\Lambda}$ is a random Schrödinger operator on $\ell^{2}(\Lambda^{\left(N\right)})$. For a fixed $N\in\mathbb{N}$, $H_{N}^{\Lambda}$ satisfies all the hypothesis of the operators studied on [3] except that it is a Schrödinger operator on $\ell^{2}(\Lambda^{\left(N\right)})$, not on $\ell^{2}(\Lambda^{N})$. This does not affect the analysis in [3], and all the results of [3] hold for $H_{N}^{\Lambda}$ for a fixed $N$.
@@
We now apply [3, Theorem 4.1] to $H_{N}^{\Lambda}$ for all $N\in\mathbb{N}$. We obtain the conclusions of the theorem for $H_{N}^{\mathbb{Z}}$ for all $N\in\mathbb{N}$, with the constants independent of $N$ unless explicitly stated. It follows that the theorem holds as stated. ∎